OpenStudy (anonymous):

Solve 8.1x – 3.2 = 7.6x – 2.4. Answer x = 0.4 x = 1.6 This is an identity and is true for all real numbers. There is no solution.

OpenStudy (anonymous):

Your goal is to get x all by itself on one side of the equation. Since we have some subtraction on both sides we'll want to add something to both sides. We'll start by adding 3.2 to both sides to get this: \[8.1x=7.6x+.8\]From here, we want to subtract 7.6x from both sides. Can you take it from here?

OpenStudy (anonymous):

8.1x-7.6x=-2.4+3.2 =>0.5x=0.8 =>x=0.8/0.5 =>x=1.6